# Implicit Neural Representations with Fourier Kolmogorov-Arnold Networks

### v2

@@ Abstract
Implicit neural representations (INRs) use neural networks to provide continuous and resolution-independent representations of complex signals with a small number of parameters. However, existing INR models often fail to capture important frequency components specific to each task. To address this issue, in this paper, we propose a Fourier Kolmogorov–Arnold network (FKAN) for INRs. The proposed FKAN utilizes learnable activation functions modeled as Fourier series in the first layer to effectively control and learn the task-specific frequency components. In addition, the activation functions with learnable Fourier coefficients improve the ability of the network to capture complex patterns and details, which is beneficial for high-resolution and high-dimensional data. Experimental results show that our proposed FKAN model outperforms three state-of-the-art baseline schemes, and improves the peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) for the image representation task and intersection over union (IoU) for the 3D occupancy volume representation task, respectively.
1 Introduction
@@
Implicit neural representations (INRs), which model continuous functions from discrete data, have gained attention for their effectiveness in representing 2D images, 3D shapes, neural radiance fields, and other complex structures [1, 2, 3, 4]. Unlike traditional convolutional neural networks (CNNs) which are limited to 3D inputs, coordinate networks use 1D vectors, providing a flexible framework for solving inverse problems in any dimension. INR models build on the multi-layer perceptron (MLP) structure and alternate between linear layers and non-linear activation functions, benefiting from its continuity nature and expressive power. MLP-based INR models avoid the locality bias problem that often restricts the effectiveness of CNNs. However, ReLU-based MLPs in coordinate networks exhibit spectral bias, prioritizing low-frequency signals. As a result, these networks learn high-frequency components more slowly [5, 6, 4, 7]. This suggests that MLPs generally capture basic patterns in real-world data, focusing on the low-frequency aspects of the target function [6, 8].
@@
To overcome the challenge of capturing high-frequency components, several approaches have been explored. Spatial encoding techniques like frequency decomposition, high-pass filtering, and Fourier features [9] help emphasize high-frequency components, while architectural modifications such as multi-scale representations [10] capture both low-frequency and high-frequency details. Additionally, methods like SIREN [2] and WIRE [11] use periodic activation functions, such as sine functions, for automatic frequency tuning [12, 11]. However, the aforementioned approaches introduce new challenges. The effectiveness of the SIREN model relies heavily on the proper selection of hyperparameters, like frequency. It is sensitive to initialization and requires careful design to prevent random variations. Moreover, due to the unknown frequency distribution of the signal, spatial encoding techniques face a mismatch between the predefined frequency bases and the signal’s inherent properties, causing an incomplete or inaccurate representation [13, 14].
@@
To address the aforementioned issues, in this paper, we propose a novel approach that enhances the hierarchical representation of INRs for improved signal reconstruction in tasks like image representations and 3D structure modeling. We develop an adaptive mapping function that can manage non-linearity and intricate frequency distributions. We hypothesize that a polynomial approximation of activation functions in the initial layer can capture fine-grained, high-frequency details [15]. Inspired by Kolmogorov–Arnold networks [16, 17], we introduce the Fourier Kolmogorov–Arnold network (FKAN) to learn task-specific frequency components for INRs. Our key contributions are summarized as follows:
@@ 2 Problem Formulation
where $\mathcal{L}$ denotes a pre-defined loss function and $N$ represents the number of training samples. In this paper, we consider the L2 loss function, i.e., $\mathcal{L}=||f\left(\boldsymbol{x}_{n};\boldsymbol{\Phi}\right)-\boldsymbol{y}_{n}||^{2}$. Also, $\boldsymbol{x}_{n}$ and $\boldsymbol{y}_{n}$ denote the input and output signal for the $n=\{1,\dots,N\}$ training sample, respectively.
Fig. 1: Illustration of the proposed FKAN model. The proposed architecture includes an FKAN block for capturing task-specific frequency components with learnable activation functions and includes $L$ hidden layers to learn non-linear patterns in the signals.
3 Proposed Fourier Kolmogorov-Arnold Network
To capture task-specific frequency components in a fine-grained manner, we propose the FKAN. Motivated by the Kolmogorov-Arnold representation theorem [19] and KANs [16], which employ learnable activation functions on edges instead of nodes as in vanilla MLPs, our proposed FKAN utilizes learnable activation functions modeled as Fourier series. This approach allows for learning a higher spectral resolution for signals. The first layer of the proposed spectral FKAN can be expressed as follows:
@@ 3 Proposed Fourier Kolmogorov-Arnold Network
where $a_{k}$ and $b_{k}$ denote the learnable Fourier coefficients, and $K$ is the number of frequency components (or grid size), which can be fine-tuned as a hyper-parameter. The proposed architecture can control and capture a wide range of frequency components, leveraging the spectral properties of the Fourier series to efficiently represent both low and high-frequency components of the input signal. Moreover, Fourier series representation has a lower training complexity compared to spline functions. The FKAN with a single layer of learnable activation functions is sufficient to achieve a high-quality spectral representation of the input signal.
Fig. 2: Comparison of the image representation between proposed FKAN and baselines.
To learn the intrinsic non-linear patterns in data, as shown in Fig. 1, we utilize $L$ hidden layers, each performing a linear transformation followed by a fixed non-linear activation function. The final layer then applies a linear transformation to generate the output signal. The non-linear activation in hidden layers plays an important role in improving the representation capacity of INRs [11]. To this end, we use the $\tanh(\cdot)$ activation function for the hidden layers. The architecture of the hidden layers is as follows:
@@
where $\boldsymbol{z}_{i}$ is the input to the $i$-th hidden layer, with $\boldsymbol{z}_{1}=\boldsymbol{\Psi}(\boldsymbol{x})$, $\boldsymbol{W}_{i}\in\mathbb{R}^{H_{i}\times H_{i+1}}$ and $\boldsymbol{b}_{i}\in\mathbb{R}^{H_{i+1}}$ are the learnable weights for the linear transformation in the $i$-th hidden layer. $\omega_{0}\in\mathbb{R}^{+}$ is a pre-defined positive scalar to control the frequency and convergence of the model, with $\omega_{0}=30$ in our implementations. In addition, we initialize the weights in the hidden layers using uniform distribution $\boldsymbol{W}_{i}\sim\mathcal{U}\left(-\sqrt{{6}/{d_{i}}},\sqrt{{6}/{d_{i}}}\right)$.
@@
where $\boldsymbol{W}_{f}\in\mathbb{C}^{H_{L}\times d_{o}}$ and $\boldsymbol{b}_{f}\in\mathbb{C}^{d_{o}}$ are the learnable weights for the linear transformation in the final layer. [^1] As mentioned in Section 2, the output signal can contain complex values. Therefore, we initialize the weights in the final layer as complex numbers to generate a complex-valued output signal. Complex-valued operations are managed based on Wirtinger calculus [21, 22]. For cases where the output signal is real-valued, the weights in the final layer are initialized as real numbers.
Table 1: Comparison of the number of parameters and performance for image representation task between methods.
Methods #Params (K) PSNR (dB) SSIM
@@
Fig. 3: Illustration of the convergence rates of the models for the image representation task.
Fig. 4: Comparison of the occupancy volume representation between proposed FKAN and baselines.
Fig. 5: Illustration of the convergence rates of the models for the occupancy volume representation task.
4 Performance Evaluation
Implementation Details: We evaluate the effectiveness of our proposed FKAN over image representation and occupancy volume representation tasks. Our experiments are conducted on an Nvidia RTX 4070 GPU with 12GB of memory. To implement the neural networks, we use PyTorch library [23] and Adam optimizer [24]. We choose $H_{1}=128$ for the latent dimension of the FKAN block with grid size $K=250$. We choose $L=4$ for the number of hidden layers with $256$, $256$, $256$, and $512$ hidden neurons in each layer, respectively. The initial learning rate is set to $0.0001$. We consider $500$ and $200$ epochs to train the models for image representation and occupancy volume representation tasks, respectively. We compare the performance of our proposed FKAN with the following baselines: SIREN [2], WIRE [11], INCODE [18], and FFN [9].
4.1 Image Representation
@@
4.2 Occupancy Volume Representation
we conduct experiments on the Thai statue dataset from the Stanford 3D Scanning Repository with WIRE system setting [11], which maps 3D coordinates (i.e., $d_{i}=3$) to signed distance function (SDF) values (i.e., $d_{o}=1$). We create an occupancy volume through point sampling on a $512\times 512\times 512$ grid. To evaluate the performance of our proposed FKAN for the occupancy volume representation task, we consider the intersection over union (IoU) metric. We plot the reconstructed 3D shapes in Fig. 4. We observe that our proposed FKAN model outperforms all the baselines. In particular, the proposed FKAN provides $0.96\%$ improvements on the IoU metric compared to the INCODE. FKAN utilizes learnable activation functions that can capture both low-frequency smooth regions and high-frequency details, resulting in the highest IoU scores. In Fig. 5, we plot the convergence rate of the models for the occupancy volume representation task. We observe that the proposed FKAN has a faster convergence compared to all baselines.
5 Conclusion
In this paper, we proposed FKAN for implicit neural signal representations. The proposed FKAN utilizes learnable activation functions modeled as Fourier series to capture task-specific frequency components and learn complex patterns of high-dimensional signals in a fine-grained manner. We investigated the performance of our proposed FKAN on two signal representation tasks, namely image representation and 3D occupancy volume representation. Experimental results demonstrate that our proposed FKAN outperforms three state-of-the-art baselines with faster convergence. It improves the PSNR and SSIM for the image representation task and IoU for the 3D occupancy volume representation task. For future work, we will consider image-denoising, image super-resolution, and neural radiance field tasks.

### v3

@@ Abstract
Implicit neural representations (INRs) use neural networks to provide continuous and resolution-independent representations of complex signals with a small number of parameters. However, existing INR models often fail to capture important frequency components specific to each task. To address this issue, in this paper, we propose a Fourier Kolmogorov–Arnold network (FKAN) for INRs. The proposed FKAN utilizes learnable activation functions modeled as Fourier series in the first layer to effectively control and learn the task-specific frequency components. The activation functions with learnable Fourier coefficients improve the ability of the network to capture complex patterns and details, which is beneficial for high-resolution and high-dimensional data. Experimental results show that our proposed FKAN model outperforms four state-of-the-art baseline schemes, and improves the peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) for the image representation task and intersection over union (IoU) for the 3D occupancy volume representation task, respectively. The code is available at github.com/Ali-Meh619/FKAN.
I Introduction
@@
Implicit neural representations (INRs), which model continuous functions from discrete data, have gained attention for their effectiveness in representing 2D images, 3D shapes, neural radiance fields, and other complex structures [1, 2, 3, 4]. Unlike traditional convolutional neural networks (CNNs) which are limited to 3D inputs, coordinate networks use 1D vectors, providing a flexible framework for solving inverse problems in any dimension. INR models build on the multi-layer perceptron (MLP) structure and alternate between linear layers and non-linear activation functions, benefiting from its continuous nature and expressive power. MLP-based INR models avoid the locality bias problem that often restricts the effectiveness of CNNs. However, rectified linear unit (ReLU)-based MLPs in coordinate networks exhibit spectral bias, prioritizing low-frequency signals. As a result, these networks learn high-frequency components more slowly [5, 6, 4, 7]. This suggests that MLPs generally capture basic patterns in real-world data, focusing on the low-frequency aspects of the target function [6, 8].
@@
To overcome the challenge of capturing high-frequency components, several approaches have been explored. Spatial encoding techniques like frequency decomposition, high-pass filtering, and Fourier features [9] help emphasize high-frequency components, while architectural modifications such as multi-scale representations [10] can capture both low-frequency and high-frequency details. Additionally, methods such as SIREN [2] and WIRE [11] use periodic activation functions, such as sine functions, for automatic frequency tuning [12, 11]. However, the aforementioned approaches introduce new challenges. The effectiveness of the SIREN model relies heavily on the proper selection of hyperparameters, like frequency. It is sensitive to initialization and requires careful design to prevent random variations. Moreover, due to the unknown frequency distribution of the signal, spatial encoding techniques face a mismatch between the predefined frequency bases and the signal’s inherent properties, causing an incomplete or inaccurate representation [13, 14].
@@
To address the aforementioned issues, in this paper, we propose a novel approach that enhances the hierarchical representation of INRs for improved signal reconstruction in tasks like image representations and 3D structure modeling. We develop an adaptive mapping function that can manage non-linearity and intricate frequency distributions. We hypothesize that a polynomial approximation of activation functions in the initial layer can capture fine-grained high-frequency details [15]. Inspired by Kolmogorov–Arnold networks (KANs) [16, 17], we introduce Fourier Kolmogorov–Arnold network (FKAN) to learn task-specific frequency components for INRs. Our key contributions are summarized as follows:
@@ II Problem Formulation
where $\mathcal{L}$ denotes a pre-defined loss function and $N$ represents the number of training samples. In this paper, we consider the $l_{2}$ loss function, i.e., $\mathcal{L}=||f\left(\boldsymbol{x}_{n};\boldsymbol{\Phi}\right)-\boldsymbol{y}_{n}||^{2}$. In addition, $\boldsymbol{x}_{n}$ and $\boldsymbol{y}_{n}$ denote the input and output signal for the $n\in\{1,\dots,N\}$ training sample, respectively.
Figure 1: Illustration of the proposed FKAN model. The proposed architecture includes an FKAN block for capturing task-specific frequency components with learnable activation functions and includes $L$ hidden layers to learn non-linear patterns in the input signals.
III Proposed Fourier Kolmogorov-Arnold Network
To capture the task-specific frequency components in a fine-grained manner, we propose the FKAN. Motivated by the Kolmogorov-Arnold representation theorem [19] and KANs [16], which employ learnable activation functions on edges instead of nodes as in vanilla MLPs, our proposed FKAN utilizes learnable activation functions modeled as Fourier series. This approach allows for learning a higher spectral resolution of signals. The first layer of the proposed spectral FKAN can be expressed as follows:
@@ III Proposed Fourier Kolmogorov-Arnold Network
where $a_{k}$ and $b_{k}$ denote the learnable Fourier coefficients, and $K$ is the number of frequency components (or grid size), which can be fine-tuned as a hyper-parameter. The proposed architecture can control and capture a wide range of frequency components, leveraging the spectral properties of the Fourier series to efficiently represent both the low-frequency and high-frequency components of the input signal. Moreover, Fourier series representation has a lower training complexity compared to spline functions. The FKAN with a single layer of learnable activation functions is sufficient to achieve a high-quality spectral representation of the input signal.
Figure 2: Comparison of the image representation between proposed FKAN and baselines.
To learn the intrinsic non-linear patterns in the data, as shown in Fig. 1, we utilize $L$ hidden layers, each performing a linear transformation followed by a fixed non-linear activation function. The final layer then applies a linear transformation to generate the output signal. The non-linear activation in hidden layers plays an important role in improving the representation capacity of INRs [11]. To this end, we use the $\tanh(\cdot)$ activation function for the hidden layers. The architecture of the hidden layers is as follows:
@@
where $\boldsymbol{z}_{i}$ is the input to the $i$-th hidden layer, with $\boldsymbol{z}_{1}=\boldsymbol{\Psi}(\boldsymbol{x})$, $\boldsymbol{W}_{i}\in\mathbb{R}^{H_{i}\times H_{i+1}}$ and $\boldsymbol{b}_{i}\in\mathbb{R}^{H_{i+1}}$ are the learnable weights for the linear transformation in the $i$-th hidden layer. $\omega_{0}\in\mathbb{R}^{+}$ is a pre-defined positive scalar to control the frequency and convergence of the model, with $\omega_{0}=30$ in our implementations. In addition, we initialize the weights in the hidden layers using uniform distribution $\boldsymbol{W}_{i}\sim\mathcal{U}(-\sqrt{{6}/{d_{i}}},\sqrt{{6}/{d_{i}}})$.
@@
where $\boldsymbol{W}_{f}\in\mathbb{C}^{H_{L}\times d_{o}}$ and $\boldsymbol{b}_{f}\in\mathbb{C}^{d_{o}}$ are the learnable weights for the linear transformation in the final layer. [^1] As mentioned in Section II, the output signal can contain complex values. Therefore, we initialize the weights in the final layer as complex numbers to generate a complex-valued output signal. Complex-valued operations are managed based on the Wirtinger calculus [21, 22]. For cases where the output signal is real-valued, the weights in the final layer are initialized as real numbers.
TABLE I: Comparison of the number of parameters and performance for image representation task between methods.
Methods #Parameters PSNR (dB) SSIM
@@
Figure 3: Illustration of the convergence rates of the models for the image representation task.
Figure 4: Comparison of the occupancy volume representation between proposed FKAN and baselines.
Figure 5: Illustration of the convergence rates of the models for the occupancy volume representation task.
IV Performance Evaluation
Implementation Details: We evaluate the effectiveness of our proposed FKAN over image representation and occupancy volume representation tasks. Our experiments are conducted on an Nvidia RTX 4070 GPU with 12GB of memory. To implement the neural networks, we use PyTorch library [23] and Adam optimizer [24]. We choose $H_{1}=128$ for the latent dimension of the FKAN block with grid size $K=270$. We choose $L=4$ for the number of hidden layers with $256$, $256$, $256$, and $512$ hidden neurons in each layer, respectively. The initial learning rate is set to $0.0001$. We consider $500$ and $200$ epochs to train the models for image representation and occupancy volume representation tasks, respectively. We compare the performance of our proposed FKAN with the following baselines: SIREN [2], WIRE [11], INCODE [18], and FFN [9].
IV-A Image Representation
@@
IV-B Occupancy Volume Representation
We conduct experiments on the Thai statue dataset from the Stanford 3D Scanning Repository with WIRE system setting [11], which maps 3D coordinates (i.e., $d_{i}=3$) to signed distance function (SDF) values (i.e., $d_{o}=1$). We create an occupancy volume through point sampling on a $512\times 512\times 512$ grid. To evaluate the performance of our proposed FKAN for the occupancy volume representation task, we consider the intersection over union (IoU) metric. We plot the reconstructed 3D shapes in Fig. 4. We observe that our proposed FKAN model outperforms all the baselines. In particular, the proposed FKAN provides $0.96\%$ improvements on the IoU metric compared to the INCODE as the second-best model. FKAN utilizes learnable activation functions that can capture both low-frequency smooth regions and high-frequency details, resulting in the highest IoU scores. In Fig. 5, we plot the convergence rate of the models for the occupancy volume representation task. We observe that the proposed FKAN has a faster convergence compared to all the baselines.
V Conclusion
In this paper, we proposed FKAN for implicit neural signal representations. The proposed FKAN utilizes learnable activation functions modeled as Fourier series to capture task-specific frequency components and learn complex patterns of high-dimensional signals in a fine-grained manner. We investigated the performance of our proposed FKAN on two signal representation tasks, namely image representation and 3D occupancy volume representation. Experimental results demonstrated that our proposed FKAN outperforms four state-of-the-art baselines with faster convergence. It improves the PSNR and SSIM for the image representation task and IoU for the 3D occupancy volume representation task. For future work, we will consider neural radiance field task.
VI Acknowledgments
This work was supported by the Canadian Foundation for Innovation-John R. Evans Leaders Fund (CFI-JELF) program grant number 42816. We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) Discovery Grant RGPIN-2023-03575 and Mitacs through the Mitacs Accelerate program grant number AWD-024298-IT33280.